# CATD: Unified Representation Learning for EEG-to-fMRI Cross-Modal Generation

### v1

@@ Abstract
Multi-modal neuroimaging analysis is crucial for a comprehensive understanding of brain function and pathology, as it allows for the integration of different imaging techniques, thus overcoming the limitations of individual modalities. However, the high costs and limited availability of certain modalities pose significant challenges. To address these issues, this paper proposed the Condition-Aligned Temporal Diffusion (CATD) framework for end-to-end cross-modal synthesis of neuroimaging, enabling the generation of functional magnetic resonance imaging (fMRI)-detected Blood Oxygen Level Dependent (BOLD) signals from more accessible Electroencephalography (EEG) signals. By constructing Conditionally Aligned Block (CAB), heterogeneous neuroimages are aligned into a potential space, achieving a unified representation that provides the foundation for cross-modal transformation in neuroimaging. The combination with the constructed Dynamic Time-Frequency Segmentation (DTFS) module also enables the use of EEG signals to improve the temporal resolution of BOLD signals, thus augmenting the capture of the dynamic details of the brain. Experimental validation demonstrated the effectiveness of the framework in improving the accuracy of neural activity prediction, identifying abnormal brain regions, and enhancing the temporal resolution of BOLD signals. The proposed framework establishes a new paradigm for cross-modal synthesis of neuroimaging by unifying heterogeneous neuroimaging data into a potential representation space, showing promise in medical applications such as improving Parkinson’s disease prediction and identifying abnormal brain regions.
@@ 1 Introduction
The BOLD signal, measured by fMRI, provides a detailed and precise mapping of brain activity[1]. The sensitivity of this signal to changes in oxygenation and deoxygenation levels in the blood provides a dynamic image of brain function. This helps to understand and track various brain diseases and is even considered the gold standard of modern functional neuroimaging[2]. Notably, BOLD fMRI has provided important insights into the brain dynamics of disorders such as ischemic stroke[3], schizophrenia[4], Alzheimer’s disease[5], focal epilepsy[6], and depression[7]. This highlights its indispensable role in diagnosing and monitoring these diseases.
@@
Although BOLD fMRI is highly regarded for its detailed imaging capabilities, acquiring such scans is not only costly[8] and time-consuming[9], but also has limitations in a variety of clinical setting applications[10, 11]. At the same time, the inherent delay between neuronal activity and haemodynamic responses captured by BOLD fMRI can limit its effectiveness in real-time brain activity monitoring[12]. In contrast, EEG has several advantages in reflecting brain activity. Because EEG directly measures brain electrical activity, it provides real-time insight with high temporal resolution, which is critical for capturing rapid dynamic changes in neuronal circuits[13]. However, the lower spatial resolution of EEG limits its ability to accurately localise activity in brain regions[14], and Fig.1 outlines the respective strengths and limitations of fMRI and EEG.
Figure 1: Comparison of the respective advantages and disadvantages of BOLD fMRI and EEG
@@
In recent years, cross-modal neuroimage synthesis has gradually become a research hotspot with the rapid development of artificial intelligence technology[15], especially with the wide application of generative models in different fields. Generative AI models, such as transformer[16], diffusion models[17, 18, 19] and generative adversarial networks (GANs)[20, 21, 22], demonstrate great potential in cross-modal data synthesis[23, 24]. For example, Yan et al. used a deep convolutional generative adversarial network (DCGAN) to reconstruct missing BOLD signals for individual participants[25]. The success of generative AI is based on the existence of some of the same underlying information and potentially relevant features in different modal information. Although EEG and fMRI are acquired in different ways, they both reflect brain activity and EEG has the advantage of low cost and fewer limitations on its use. It has been found that there is a strong correlation between microstate transitions in EEG signals and BOLD signals[26] and that BOLD functional connectivity correlates with functional connectivity of EEG activity[27]. By simultaneously recording EEG and fMRI signals, researchers can observe the relationship between EEG activity and metabolic activity in the brain. When electrical brain activity increases, blood oxygen levels in the brain also increase, suggesting a one-to-one correspondence between changes in electrical brain activity and the state of metabolic activity in the brain[28, 29]. This correlation provides theoretical possibilities for cross-modal reconstruction and temporal resolution enhancement of fMRI using generative AI and EEG.
@@
EEG and BOLD signals are highly heterogeneous time-series data, requiring powerful generative AI models for cross-modal synthesis. Currently, diffusion transformer model[30] has demonstrated excellent performance in several areas such as image generation. The emergence of Sora[31] further demonstrates the potential application of diffusion models in this direction of temporal data processing. Drawing inspiration from these advancements, the CATD framework for the unified representation of EEG and BOLD signals is proposed. This framework is the first to achieve cross-modal synthesis of high-dimensional, heterogeneous brain functional data using a diffusion model. The novelties and contributions of this paper can be summarized in the following points:
@@
    A new paradigm based on generative AI for unified representation of neuroimaging is proposed. As far as we know, it is the first time a diffusion-driven end-to-end framework is developed for EEG-to-fMRI synthesis. Utilizing low-cost, accessible EEG signals, the proposed CATD framework is capable to synthesis high-cost, difficult-to-access BOLD signals. The proposed framework enables high-quality and stable cross-modal generation from EEG to BOLD signals, bridging the gap between different neuroimaging modalities.
@@
    The CAB module is designed to align high-temporal, low-spatial resolution EEG signals with low-temporal, high-spatial resolution BOLD signals within a latent space, facilitating a unified representation across modalities. The combination with DTFS module also leverages EEG’s superior temporal resolution to improve the temporal super-resolution of BOLD signals, capturing detailed brain dynamics that surpass traditional methods.
@@ 2.1 Overview
BOLD signals are valuable in reflecting brain activity and are essential for analysing, diagnosing and treating brain disorders. However, BOLD signal acquisition is not possible in some patients due to medical conditions or other limitations. In response, the paper proposed the CATD framework, an innovative EEG-to-BOLD signal conversion model based on Scalable Diffusion Models with Transformers (DiT)[30]. This approach addresses the limitations of cross-modal reconstruction and temporal super-resolution of whole-brain BOLD signals, which are not possible with existing techniques. The CATD framework addresses the challenges posed by high dimensionality and asymmetry between EEG and BOLD data through a novel heterogeneous alignment method that facilitates dimensional matching and enhances signal compatibility. At the same time, it employs a cross-attention mechanism to efficiently generate cross-modal EEG-modulated BOLD signals. The proposed DTFS achieves the control of the EEG sampling rate by sliding the sampler, which in turn achieves the enhancement of the temporal resolution of the output BOLD signal. Fig.2 depicts the complete structure and functionality of the CATD framework.
Figure 2: The overall framework of the proposed CATD. The upper part of the figure shows how two different dimensions of data are processed to achieve the initial alignment. The lower half shows the reconstruction pipeline of the BOLD signal under the control of the EEG condition based on the DiT structure.
@@ 2.2 Data Alignment Method
EEG and BOLD signals are high-dimensional signals, one with high temporal resolution and the other with high spatial resolution, highly heterogeneous in scale, and both need to be processed to the same scale using the following methods to perform integration operations. For preprocessing BOLD signals, we followed the method by Yan et al. [25], utilizing freesurfer[32] software to register brain BOLD activity to the cerebral cortex. The data is then aligned using the fsaverage template and downsampled to the fsaverage4 template, creating a functional map of cortical BOLD activity. These signals are fed into the encoder of a pretrained Variational Autoencoder (VAE) [33] for dimensionality reduction. The reduced data is segmented into patches and transformed into input tokens via an embedding layer, producing the initial state $x_{0}$ for the diffusion model training.
EEG signals, corresponding to the first 6 seconds of each BOLD functional map, are selected due to the 6-second delay between neuronal activity and blood oxygenation response [34, 35]. Feature extraction and dimensionality reduction are performed using a dynamic temporal-frequency analysis method based on the short-time Fourier transform. The reduced EEG data is then divided into segments and converted into tokens that match the BOLD frames in dimension and number through the EEG embedding layer. These tokens serve as the conditional signals $c$ for the diffusion model, enabling alignment of the high-dimensional, heterogeneous time-series EEG data with BOLD signals.
For the BOLD signal super-resolution task, the sample rate of the EEG signal is controlled using the DTFS module. During training, overlapping EEG signal samples are employed. Specifically, the EEG signal is sampled at one-third of its original interval while maintaining the same overall sampling duration. This fine-grained temporal segmentation approach allows the reconstructed BOLD signal to achieve three times the temporal resolution of the original signal. As a result, this method generates more refined BOLD signal sequences, significantly enhancing temporal resolution and providing a more detailed representation of brain activity.
@@ 2.3 EEG to BOLD Diffusion
Decode Reconstructed Signal:
@@
  // Decode reconstructed BOLD signal
@@ 2.3.1 Basic ideas
As one of the highest performing generative AI models available, diffusion models are known for their stable training process and superior quality of generated output. These models work by gradually transforming the data distribution into a Gaussian distribution through a forward process, and then learning the reverse transformation to generate new data samples. In the forward phase, the diffusion model starts with the original data $x_{0}$ and gradually increases the noise over a number of time steps, eventually reaching an almost completely random state $x_{T}$, a process that can be described by the following Gaussian process:
@@
where $\bar{\alpha}_{t}$ is a hyperparameter that decreases with time. This process is characterised by a Markov chain and each step is usually controlled by a variance preserving transformation.
The inverse process aims to reconstruct the original data from the noisy state. This is done by training a neural network to predict the noise added at each step of the forward process, and then iteratively removing this noise to recover the original data $x_{0}$. This inverse process can be described by the following equation:
@@
where $\mu_{\theta}\left(x_{t},c\right)$ and $\Sigma_{\theta}\left(x_{t},c\right)$ are predicted by a neural network with parameter $\theta$. We have adopted a transformer architecture for this network because of its flexibility and excellent ability to capture long-range dependencies, which makes it particularly suitable for processing time-series signals such as EEG. The inherent attentional mechanism of the transformer allows for the integration of conditional information denoted by $c$ through a cross-attentional mechanism. This integration guides the generation process, thus enhancing the applicability of the model to cross-modal generation tasks involving high-dimensional, heterogeneous data.
@@ 2.3.2 Architectures
As shown in Fig.2, our proposed EEG to BOLD diffusion model consists of a prediction network that consists of CAB, several layers of Transformer Blocks connected in series, an input embedding layer, and an output section. Each Transformer Block contains mainly one layer of multi-head self-attention mechanism and one layer of multi-head cross-attention mechanism. Among them, the cross-attention layer is responsible for incorporating the condition information into the network. The CAB integrates the EEG markers as condition $c$ as well as the diffusion step $t,t\in\left\{0,1,\dots,T\right\}$, and delivers this integrated conditional information to the cross-attention network of each Transformer Module as a way to achieve an effective input of conditional information.
@@ 2.3.3 Loss Function
In discussing the training process of the diffusion model, it is crucial to consider the consistency of the entire reconstruction process, which requires the model to complete the computation of all time steps $T$ before each parameter update. While this approach ensures the comprehensiveness of the learning process, it has a significant negative impact on the convergence speed, stability and optimisation efficiency of the model. At the same time, it also places a high demand on computational resources. In view of this, we decided to simplify Eq.3. Instead of seeking to minimise the KL scatter in terms of distributional similarity across all time steps, we instead turned to minimising the prediction error in each time step. This approach not only simplifies the computational process but also helps to improve the training efficiency and stability of the model. The expression of the simplified loss function is as follows:
@@ 3.1 Experiment Settings
Figure 3: (a) Radar plot of the real BOLD signal, synthetic BOLD signal, real + synthetic BOLD signal, and the results of the ablation of the conditioned blocks in the motion imagery state and the resting state classification results. (b) Radar plot of different frequency bands of EEG synthesized BOLD signals in the motion and resting states after our proposed CATD framework. (c) Radar plot of real EEG signals synthesized BOLD signals, and the combination of the two for the prediction of Parkinson’s disease in a medical decision support experiment.

### v3

@@ Abstract
Multi-modal neuroimaging analysis is crucial for a comprehensive understanding of brain function and pathology, as it allows for the integration of different imaging techniques, thus overcoming the limitations of individual modalities. However, the high costs and limited availability of certain modalities pose significant challenges. To address these issues, this paper proposes the Condition-Aligned Temporal Diffusion (CATD) framework for end-to-end cross-modal synthesis of neuroimaging, enabling the generation of functional magnetic resonance imaging (fMRI)-detected Blood Oxygen Level Dependent (BOLD) signals from more accessible Electroencephalography (EEG) signals. By constructing Conditionally Aligned Block (CAB), heterogeneous neuroimages are aligned into a latent space, achieving a unified representation that provides the foundation for cross-modal transformation in neuroimaging. The combination with the constructed Dynamic Time-Frequency Segmentation (DTFS) module also enables the use of EEG signals to improve the temporal resolution of BOLD signals, thus augmenting the capture of the dynamic details of the brain. Experimental validation demonstrates that the framework improves the accuracy of brain activity state prediction by 9.13% (reaching 69.8%), enhances the diagnostic accuracy of brain disorders by 4.10% (reaching 99.55%), effectively identifies abnormal brain regions, enhancing the temporal resolution of BOLD signals. The proposed framework establishes a new paradigm for cross-modal synthesis of neuroimaging by unifying heterogeneous neuroimaging data into a latent representation space, showing promise in medical applications such as improving Parkinson’s disease prediction and identifying abnormal brain regions.
@@ 1 Introduction
The BOLD signal, measured by fMRI, provides a detailed and precise mapping of brain activity[1]. The sensitivity of this signal to changes in oxygenation and deoxygenation levels in the blood provides a dynamic image of brain function. This helps to understand and track various brain diseases and is considered the gold standard of modern functional neuroimaging[2]. Notably, BOLD fMRI has provided important insights into the brain dynamics of disorders such as ischemic stroke[3], schizophrenia[4], Alzheimer’s disease[5], focal epilepsy[6], and depression[7]. This highlights its indispensable role in diagnosing and monitoring these diseases.
Figure 1: Comparison of Advantages and Disadvantages of BOLD fMRI and EEG
@@
Although BOLD fMRI is highly regarded for its detailed imaging capabilities, acquiring such scans is not only costly[8] and time-consuming[9] but also subject to limitations in various clinical applications[10, 11]. Specifically, its poor temporal resolution limits the ability to accurately capture rapid neural dynamics that occur within milliseconds. The hemodynamic response delay inherent to fMRI makes it less suitable for real-time applications, such as brain-machine interfaces or neurofeedback systems[12]. At the same time, the high costs and limited portability of fMRI make it inaccessible for many patients, particularly in low-resource settings. In contrast, EEG has several advantages in reflecting brain activity. Because EEG directly measures brain electrical activity, it provides real-time insight with high temporal resolution, which is critical for capturing rapid dynamic changes in neuronal circuits[13]. However, EEG’s low spatial resolution significantly hinders its ability to localize brain activity accurately, limiting its use in tasks requiring precise spatial mapping[14]. The low signal-to-noise ratio of EEG data and its susceptibility to artifacts (e.g., muscle movement, eye blinks) also pose challenges for reliable interpretation, especially in complex cognitive or clinical scenarios. Fig.1 outlines the respective strengths and limitations of fMRI and EEG.
@@
In recent years, with the advancements in information technology[15, 16] and artificial intelligence techniques[17, 18, 19, 20], cross-modal neuroimage synthesis has gradually become an active area of research, especially with the wide application of generative models in different fields. Generative AI models, such as transformer[21], diffusion models[22, 23, 24] and generative adversarial networks (GANs)[25, 26, 27], demonstrate great potential in cross-modal data synthesis[28, 29], and their applications in brain imaging[30] and brain function[31, 32] are also gradually increasing. For example, Yan et al. used a deep convolutional generative adversarial network (DCGAN) to reconstruct missing BOLD signals for individual participants[33]. Calhas et al.[34] conducted early research on generative AI for EEG-to-fMRI synthesis, proposing an Autoencoder(AE)-based model for this purpose. In our experiments, we compared our proposed method to their AE-based model and demonstrated significant performance improvements in cross-modal synthesis. The success of generative AI is based on the existence of some of the same underlying information and potentially relevant features in different modalities. Although EEG and fMRI are acquired in different ways, they both reflect brain activity and EEG has the advantage of low cost and fewer limitations on its use. It has been found that there is a strong correlation between microstate transitions in EEG signals and BOLD signals[35] and that BOLD functional connectivity correlates with functional connectivity of EEG activity[36]. By simultaneously recording EEG and fMRI signals, researchers can observe the relationship between EEG activity and metabolic activity in the brain. When electrical brain activity increases, blood oxygen levels in the brain also increase, suggesting a one-to-one correspondence between changes in electrical brain activity and the state of metabolic activity in the brain[37, 38]. This correlation provides a theoretical basis for cross-modal generation and temporal resolution enhancement of fMRI using generative AI and EEG.
@@
EEG and BOLD signals are highly heterogeneous time-series data, requiring powerful generative AI models for cross-modal synthesis. Currently, diffusion transformer model[39] has demonstrated excellent performance in several areas such as image generation. The emergence of Sora[40] further demonstrates the potential application of diffusion models for temporal data synthesis. Drawing inspiration from these advancements, the CATD framework for the unified representation of EEG and BOLD signals is proposed. This framework is the first to achieve cross-modal synthesis of high-dimensional, heterogeneous brain functional data using a diffusion model. The novelties and contributions of this paper can be summarized in the following points:
@@
    A new paradigm based on generative AI for unified representation of neuroimaging is proposed. As far as we know, it is the first time a diffusion-driven end-to-end framework is developed for EEG-to-fMRI synthesis. Utilizing low-cost, accessible EEG signals, the proposed CATD framework is capable of synthesizing high-cost, difficult-to-access BOLD signals. The proposed framework enables high-quality, stable cross-modal generation from EEG to BOLD signals, bridging the gap between different neuroimaging modalities.
@@
    The CAB module is designed to align high-temporal, low-spatial resolution EEG signals with low-temporal, high-spatial resolution BOLD signals within a latent space, facilitating a unified representation across modalities. The combination with DTFS module also leverages EEG’s superior temporal resolution to improve the temporal super-resolution of BOLD signals, capturing detailed brain dynamics that outperform traditional methods.
The remainder of this paper is organized as follows. Section 2 introduces the proposed CATD framework in detail, including its structural design and methodology. Section 3 describes the experimental setup, datasets used, implementation details, and the experimental results. Section 4 provides a comprehensive analysis and discussion of the experimental results as well as the reasons for their occurrence, the applicability and potential of the framework. Finally, Section 5 concludes the paper by summarizing the key contributions and highlighting future research directions.
@@ 2.1 Overview
BOLD signals are valuable in reflecting brain activity and are essential for analysing, diagnosing and treating brain disorders. However, BOLD signal acquisition is not possible in some patients due to medical conditions or other limitations. In response, the paper proposes the CATD framework, an innovative EEG-to-BOLD signal conversion model based on Scalable Diffusion Models with Transformers (DiT)[39]. This approach addresses the limitations of cross-modal generation and temporal super-resolution of whole-brain BOLD signals, which are not possible with existing techniques. The CATD framework addresses the challenges posed by high dimensionality and asymmetry between EEG and BOLD data through a novel heterogeneous alignment method that facilitates dimensional matching and enhances signal compatibility. At the same time, it employs a cross-attention mechanism to efficiently generate cross-modal EEG-modulated BOLD signals. The proposed DTFS achieves the control of the EEG sampling rate by sliding the sampler, which in turn achieves the enhancement of the temporal resolution of the output BOLD signal. Fig.2 depicts the complete structure and functionality of the CATD framework.
Figure 2: The overall framework of the proposed CATD. The upper part of the figure shows how two different dimensions of data are processed to achieve the initial alignment. The lower half shows the generation pipeline of the BOLD signal under the control of the EEG condition based on the DiT structure.
@@ 2.2 Data Alignment Method
EEG and BOLD signals are high-dimensional signals, one with high temporal resolution and the other with high spatial resolution, highly heterogeneous in scale, and both need to be processed to the same scale using the following methods to perform integration operations. For preprocessing BOLD signals, we followed the method by Yan et al. [33], utilizing freesurfer[41] software to register brain BOLD activity to the cerebral cortex. Individual structural images were processed to generate cortical surface meshes, and structural and functional images were aligned using boundary-based registration with FsFast. The fMRI data were aligned to the fsaverage template and subsequently downsampled to the fsaverage4 template, creating a functional map with 2562 vertices per hemisphere using the mri_surf2surf function in FreeSurfer. Additionally, a 6-mm full-width at half-maximum (FWHM) Gaussian smoothing kernel was applied to the fMRI data in surface space. These processed signals are fed into the encoder of a pretrained Variational Autoencoder (VAE)[42], specifically sd-vae-ft-mse from the Diffusers library [43], for dimensionality reduction. The reduced data are then segmented into patches and transformed into input tokens via an embedding layer, producing the initial state $x_{0}$ for diffusion model training.
EEG signals, corresponding to the first 6 seconds of each BOLD functional map, are selected due to the 6-second delay between neuronal activity and blood oxygenation response [44, 45]. Feature extraction and dimensionality reduction are performed using a dynamic time-frequency analysis method based on the short-time Fourier transform. The reduced EEG data are then divided into segments and converted into tokens that match the BOLD frames in dimension and number through the EEG embedding layer. These tokens serve as the conditional signals $c$ for the diffusion model, enabling alignment of the high-dimensional, heterogeneous time-series EEG data with BOLD signals.
For the BOLD signal super-resolution task, the sampling rate of the EEG signal is controlled using the DTFS module. During training, overlapping EEG signal samples are employed. Specifically, the EEG signal is sampled at one-third of its original interval while maintaining the same overall sampling duration. This fine-grained temporal segmentation approach allows the generated BOLD signal to achieve three times the temporal resolution of the original signal. As a result, this method generates more refined BOLD signal sequences, significantly enhancing temporal resolution and providing a more detailed representation of brain activity.
@@ 2.3 EEG to BOLD Diffusion
Decode Generated Signal:
@@
  // Decode generated BOLD signal
@@ 2.3.1 Basic ideas
As one of the highest performing generative AI models available, diffusion models are known for their stable training process and superior quality of generated output. These models work by gradually transforming the data distribution into a Gaussian distribution through a forward process, and then learning the reverse transformation to generate new data samples. In the forward phase, the diffusion model starts with the original data $x_{0}$ and gradually increases the noise over multiple time steps, eventually reaching an almost completely random state $x_{T}$, a process that can be described by the following Gaussian process:
@@
where $\bar{\alpha}_{t}$ is a hyperparameter that decreases over time. This process is characterised by a Markov chain and each step is typically governed by a variance-preserving transformation.
The inverse process seeks to reconstruct the original data from the noisy state. This is done by training a neural network to predict the noise added at each step of the forward process, and then iteratively removing this noise to recover the original data $x_{0}$. This inverse process can be described by the following equation:
@@
where $\mu_{\theta}\left(x_{t},c\right)$ and $\Sigma_{\theta}\left(x_{t},c\right)$ are predicted by a neural network with parameter $\theta$. We adopt a transformer architecture for this network because of its flexibility and excellent ability to capture long-range dependencies, which makes it particularly suitable for processing time-series signals such as EEG. The inherent attention mechanism of the transformer allows for the integration of conditional information denoted by $c$ through a cross-attention mechanism. This integration guides the generation process, thereby enhancing the model’s applicability to cross-modal generation tasks involving high-dimensional, heterogeneous data.
@@ 2.3.2 Architectures
As shown in Fig.2, our proposed EEG to BOLD diffusion model consists of a prediction network that consists of CAB, several layers of Transformer Blocks connected in series, an input embedding layer, and an output section. Each Transformer Block primarily contains one layer of a multi-head self-attention mechanism and one layer of multi-head cross-attention mechanism. The cross-attention layer is responsible for incorporating the conditional information into the network. The CAB integrates EEG markers as condition $c$ as well as the diffusion step $t,t\in\left\{0,1,\dots,T\right\}$, and delivers this integrated conditional information to the cross-attention network of each Transformer Module as a way to achieve an effective input of conditional information.
@@ 2.3.3 Loss Function
In discussing the training process of the diffusion model, it is crucial to consider the consistency of the entire generation process, which requires the model to complete the computation of all time steps $T$ before each parameter update. While this approach ensures the comprehensiveness of the learning process, it has a significant negative impact on the convergence speed, stability and optimisation efficiency of the model. At the same time, it also places a high demand on computational resources. In view of this, we decided to simplify Eq.3. Instead of minimizing the KL divergence to ensure distributional similarity across all time steps, we instead turned to minimising the prediction error in each time step. This approach not only simplifies the computational process but also helps to improve the training efficiency and stability of the model. The expression for the simplified loss function is as follows:
@@ 3.1 Experiment Settings
Figure 3: (a) Radar plot comparing real BOLD signals, generated BOLD signals, their combination, and ablation results of conditioned blocks for motor imagery and resting state classification. (b) Radar plot of the prediction results of motor imagery and rest state for the real BOLD signal, the data generated by our method, and the data generated by the compared AE-based method. (c) Radar plot of BOLD signals synthesized from different EEG frequency bands in motor imagery and resting states using the proposed CATD framework. (d) Radar plot of BOLD signals synthesized from real EEG signals, and their combination, for predicting Parkinson’s disease in a clinical decision support experiment.